OpenStudy (anonymous):

f(x)= 3x – 7 and g(x) = –2x – 6. Find (ƒ ∘ g)(4).

OpenStudy (mathstudent55):

Find g(4) then input that value into f(x) (fog)(4) = f(g(4)) = f(-2(4) - 6) = f(-14) = 3(-14) - 7 = -42 - 7 = -49